OpenStudy (bimbim2697):

WHAT IS THE EQUATION FOR THE LINE THAT IS PARRALEL TO THE GIVEN LINE AND THAT PASSES THROUGH A GIVEN POINT ? Y=-3X+6;(0,1)

OpenStudy (turingtest):

hey, where did you go?

OpenStudy (turingtest):

Since you left all I will leave you is the formula:\[y-y_0=m(x-x_0)\]where\[(x_0,y_0)\]is your point in the problem.

OpenStudy (bimbim2697):

HUH?

OpenStudy (turingtest):

the formula for a line is:\[y-y_0=m(x-x_0)\]where\[(x_0,y_0)\]is your point in the problem. m represents the slope. It is the number next to the x. What is m in your problem?

OpenStudy (turingtest):

what is the slope of the line\[y=-3x+6\]I just said the slope is the number next to the x...

OpenStudy (anonymous):

if a line l1 is parallel to another line l2, then l1 &l2 have same slope. so as per the line equation => y=mx +c (where m is the slope and c is the constant) so the l2 => y=3x+c to find 'c' we need to substitute the given points to it 1=3(0)+c =>c=1 so your required line is =>y=3x+1

OpenStudy (anonymous):

Since it is parallel, the slope is -3 y = -3x + b We know that (0, 1) implies that x=0 and y=1 1 = -3 + b Now solving for b we have b = 4 Therefore, y = -3x + 4

OpenStudy (anonymous):

There is a shortcut which could reduce this problem into much easier things.

OpenStudy (bimbim2697):

i GOT Y=3X+1 SO I WENT WITH DHASNI ON THIS ON

OpenStudy (bimbim2697):

E

OpenStudy (turingtest):

Hey FFM, everybody's wrong, right?

OpenStudy (anonymous):

any line can be written as y=mx+c. A line parallel to other line have the same slope as the parallel line. And c can be found out by substituting the given point in thre obtained equation say here y=3x+c. Now x=0 and y=1 gives c=1.And so the required equation is y=3x+1